# Reliability, and Global Privacy

### v3

@@ VI-2 LDP at Relay
The client $k$ computes the partial sum $\bm{W}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})$, as in (8), where $\mathcal{U}_{k}^{t}$ contains the received updates by client $k$ and its own update. Let $\bm{M}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$ be the aggregated privacy noise. Following (34), we have $\nu_{k}^{t}=\mathrm{var}(\bm{M}_{k}^{t})$. When an eavesdropper (possibly the server) observes the masked partial sum $\bm{W}_{k}^{t}$, it should not be able to distinguish the identity of any participating client, i.e., it should not confidently infer whether client $j$’s contribution is included in $\bm{W}_{k}^{t}$. The privacy comes from two stochastic factors: (i) the random participation of the neighbors of client $k$, and (ii) the aggregated privacy noise, yet the privacy noises $\bm{N}_{m}^{t}$ are correlated across clients due to the coordinated nature of the secret keys. Let $\bm{W}_{k}^{t,-j}\triangleq\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})$ denote the partial sum excluding client $j$’s contribution. Similarly, define $\bm{M}_{k}^{t,-j}=\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}$, and its variance $\nu_{k}^{t,-j}=\mathrm{var}(\bm{M}_{k}^{t,-j})$. While the network topology renders both $\nu_{k}^{t}$ and $\nu_{k}^{t,-j}$ random variables, this challenge can be effectively addressed through an appropriate choice of the Bernstein parameter [7]. Moreover, the challenge of the presence of the unbounded r.v. $\bm{N}_{j}^{t}$ in $\ell_{2}$-sensitivity can be addressed using Chi-squared tail bound. The privacy of the presence of client $j$’s contributions in $k$-th partial sum is formalized below.
Theorem VI.2 (LDP for Each Individual Contribution).
@@
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.2, then the variance of $\bm{M}_{k}^{t}$ is given by $\nu_{k}^{t,-j}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\bm{I}_{D}$, and $\bar{\nu}_{k}^{-j}=\mathbb{E}[\nu_{k}^{t,-j}]=\left(\lambda^{2}\sum_{m=1}^{K}(1-p_{k,m})g_{m,k}^{2}+\bm{\Lambda}_{k}^{\mathrm{E},-j}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{\mathrm{E},-j})^{\top}\right)$, where $\bm{\Lambda}_{k}^{\mathrm{E},-j}=\mathbb{E}[\bm{\Lambda}_{k}^{t,-j}]$. For $\forall\delta_{0}$, $r_{1}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\geq r_{1})\leq\delta^{\prime}$, [^3] The choice of Bernstein parameter $r_{1}$ is discussed in [32, Appendix F]. the partial sum $\bm{W}_{k}^{t}$ is $(\epsilon_{k,j}^{(2)},(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(2)}))$-differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any participating node $j$, where
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t,-j}\in\mathcal{S})\leq e^{\epsilon_{k,j}^{(2)}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S})+(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(2)}),$ (45)
@@
Apart from the participating identity, when the transmission from client $j$ is perturbed, i.e., when $\Delta\bm{\Theta}_{j}^{t}=\bm{u}$ is perturbed to $\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime}$, the aggregated partial sums should not fluctuate significantly. This privacy guarantee is formalized below.
Theorem VI.3 (LDP Against Perturbation of Local data).
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.2, then its variance is $\nu_{k}^{t}=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}$. The $\bar{\nu}_{k}=\mathbb{E}[\nu_{k}^{t}]$ and $\bm{\Lambda}_{k}^{\mathrm{E}}=\mathbb{E}[\bm{\Lambda}_{k}^{t}]$ are specified in Theorem VI.2. For $r_{1}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{k}-\bar{\nu}_{k}\rvert\geq r_{1})\leq\delta^{\prime}$, the partial sum $\bm{W}_{k}^{t}$ is $(\epsilon_{k,j}^{(3)},(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(3)})$-differentially private in protecting the local date of any participating node $j$ from perturbations, where
$\epsilon_{k,j}^{(3)}=\begin{cases}2\cdot\left[2\log\left(\frac{1.25}{\delta_{j}^{(3)}}\right)\right]^{\frac{1}{2}}\frac{\lvert g_{k,j}\rvert R}{\sqrt{\bar{\nu}_{k}-r_{1}}}\;\;\;\;\;\mathrm{if}\;p_{k,j}\geq 0,\\ \hskip 42.67912pt\;\;\;\;\;0\;\;\;\;\;\;\hskip 42.67912pt\;\;\;\mathrm{if}\;p_{k,j}=0.\end{cases}$ (46)
@@
$\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$ $\leq e^{\epsilon_{k,j}^{(3)}}\mathrm{P_{r}}(\bm{W}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
@@ VI-3 LDP at the Server
Suppose the server receives multiple partial sums from clients, denoted by $\mathcal{V}^{t}$, but fails to reconstruct the global model using the combinator $\bm{c}_{f^{\prime}}$ ($\bm{c}_{f^{\prime}}$-Failure). In that case, the resulting aggregation, $\bm{F}_{f^{\prime}}^{t}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t,I}+\bm{N}_{m}^{t})$, is noisy, and the effective noise in the aggregation is $\bm{N}_{F_{f^{\prime}}}^{t}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$. From $\bm{F}_{f^{\prime}}^{t}$, the server should not infer the participation of client $j$ at the relaying client $k$. The variance of $\bm{N}_{F_{f^{\prime}}}^{t}$ is $\nu_{F_{f^{\prime}}}^{t}\bm{I}_{D}=\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})\bm{I}_{D}$ (specified in Appendix I), and its expectation is $\bar{\nu}_{F_{f^{\prime}}}^{t}=\mathbb{E}[\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})]$. The stochasticity providing privacy for client $j$’s contribution are: (i) the local noise $\bm{N}_{j}^{t}$ added by client $j$; (ii) the intermittent nature of the links $\tau_{k,j}$ from client $j$ to its neighbors $k$; and (iii) the intermittent nature of the link $\tau_{k}$ from client $k$ to the server. Let $\bm{F}_{f^{\prime}}^{t,\{-j,k\}}$ be the aggregated results excluding client $j$’s contribution at relay $k$. Similarly, define $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}$ and $\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}$. The privacy is formalized below.  [^4] Note that, due to the correlation among the secret keys, the standard composition theorem of DP [3] cannot apply.
Theorem VI.4 (LDP in $\bm{c}_{f^{\prime}}$-Failure).
@@
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.2 and the combinator $\bm{c}_{f^{\prime}}$, then for any $r_{2}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{f^{\prime}}^{t,\{k,-j\}}-\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}\rvert\geq r_{2})\leq\delta^{\prime}$, [^5] The choice of Bernstein parameter $r_{2}$ is discussed in [32, Appendix F]. then, in $\bm{c}_{f^{\prime}}$-Failure, $\bm{F}_{f^{\prime}}^{t}$ is $(\epsilon_{f^{\prime},\{k,j\}}^{(4)},(1-p_{k})(1-p_{k,j})(\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime})$ differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any participating client $j$ at relay $k$, where
@@
$\epsilon_{f^{\prime},\{k,j\}}^{(4)}\hskip-2.84526pt=\hskip-2.84526pt\left[\hskip-1.42262pt2\log\left(\hskip-2.84526pt\frac{1.25}{\delta_{f^{\prime},\{k,j\}}^{(4)}}\hskip-2.84526pt\right)\hskip-4.2679pt\right]^{\frac{1}{2}}\hskip-5.69054pt\frac{\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})})}{\sqrt{\bar{\nu}_{f^{\prime}}^{t,\{k,-j\}}-r_{2}}}.$ (48)
@@
Moreover, for any $\bm{u},\bm{u^{\prime}}\in\mathbb{B}_{D}(R)$, it holds that
@@
$\epsilon_{k,j,f^{\prime}}^{(5)}=2\left[2\log\left(\frac{1.25}{\delta_{j}^{(5)}}\right)\right]^{\frac{1}{2}}\frac{\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert R}{\sqrt{\bar{\nu}^{t,f^{\prime}}-r_{2}}},$ (51a)
$\tilde{p}_{j,f^{\prime}}=\hskip-2.84526pt\sum_{\mathcal{V}_{j,f^{\prime}}^{t}\in\{\mathcal{V}_{j}\cap\mathcal{V}_{f^{\prime}}\}}\prod_{k_{1}\in\mathcal{V}_{j}^{t}}(1-p_{k_{1}})\cdot p_{k_{1},j}\hskip-2.84526pt\hskip-8.53581pt\prod_{k_{2}\in\{\mathcal{V}_{j}\cap\mathcal{V}_{f^{\prime}}\}\setminus\mathcal{V}_{j,f^{\prime}}^{t}}\hskip-8.53581ptp_{k_{2}}.$ (51b)
The $\tilde{p}_{j,f^{\prime}}$ denotes the probability of client $j$’s contribution not being conveyed by those participating in $\bm{c}_{f^{\prime}}$-combination, and $\mathcal{V}_{f^{\prime}}\triangleq\{k:c_{f^{\prime},k}\neq 0\}$.
@@
With a slight abuse of the paper organization, we examine privacy via obfuscation by other local models in successful server aggregation. Unlike standard LDP, where noise is explicitly added to each input, privacy arises naturally from the randomness or uncertainty in the other local models. Nevertheless, we can leverage results from LDP theory to derive the following results. Specifically, the server should not see significant changes in the global model $\bm{M}^{t}=\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ resulting from the presence or perturbation of $\Delta\bm{\Theta}_{j}^{t}$. To facilitate theoretical analysis, we assume $\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])=\mathrm{Diag}(\bm{I}_{D}\zeta^{2})$.
Theorem VI.5 (Privacy in Global Model).
@@
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. Moreover, for any $\bm{u},\bm{u^{\prime}}\in\mathbb{B}_{D}(r_{3})$, it holds that
@@
$\leq e^{\epsilon^{(7)}}\mathrm{P_{r}}(\bm{M}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})+\delta^{(7)}.$ (54)
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$, where
@@
Remark VI.1.
When $K\rightarrow\infty$, we observe $\epsilon^{(6)}\rightarrow 0$ and $\epsilon^{(7)}\rightarrow 0$, which indicates that secure aggregation in IID settings is not differentially private for large systems when clients have equal importance.
VII Reliability
This section evaluates the server’s ability to recover the exact global model. To achieve exact recovery, the server must receive at least KsK-s complete partial sums. Otherwise, the server aggregation is not meaningful. This leads us to the following three disjoint failure scenarios.
(i) In the gradient-sharing phase, at least $s$ clients do not receive all $s$ masked gradients from its neighbors. Then no matter how good the client-to-server channels are, the server cannot perform a meaningful aggregation. The probability of this event is given by $P_{1}$ in [31, (10)].
(ii) In the gradient-sharing phase, all transmissions are successful, but there are at least $s$ client-to-server transmission failures. As a result, although all partial sums computed at clients are complete, fewer than KsK-s of them eventually arrive at the server. The probability of this event is given by $P_{2}$ in [31, (11)].
(iii) In the gradient-sharing phase, if $v$ (fewer than $s$) clients do not receive the masked gradients from all of its $s$ neighbors, and if more than svs-v transmissions from clients holding complete partial sums to the server fail, it is impossible to recover the global model. The probability of this event is given by $P_{3}$ in [31, (14)].
For all remaining scenarios beyond the three cases above, the global model can be reconstructed accurately. In other words, SecCoGC yields a binary recovery outcome for the global model, following $\mathrm{Ber}(1-P_{O})$, where
$P_{O}=P_{1}+P_{2}+P_{3}.$ (56)
One may wonder what $P_{O}$ can influence, especially given that the global model is perfect whenever it can be recovered. If the communication repeats until the server can successfully recover the global model, $P_{O}$ decides the communication overhead per round, detailed in [31, Remark 3]. In this case, the convergence is guaranteed [33]. If the local training is performed consecutively based on its own latest trained local model when the server fails to recover the global model, $P_{O}$ will statistically affect the length of local accumulation. In this case, the convergence rate depends on $P_{O}$, as will be discussed in the next section. In both designs, the cost-efficient choice of $s$ in SecCoGC is similar to [31, Section V]
VIII Convergence
This section provides a non-convex convergence rate analysis of SecCoGC under an arbitrary choice of local solvers. The analysis also applies to the general CoGC and GC. To begin with, we present several assumptions concerning LOFs that are used throughout the analysis.
Assumption VIII.1 (Smoothness).

### v4

@@ V-B LDP at the Relaying Client
When an eavesdropper (possibly the server) observes the partial sum $\bm{S}_{k}^{t}$, it should not be able to identify any participating client $j$, i.e., it should not confidently infer whether client $j$’s contribution $g_{k,j}(\Delta\bm{\Theta}_{j}^{t}+\bm{N}_{j}^{t})$ is included in $\bm{S}_{k}^{t}$. In this case, privacy stems from: (i) the random participation of client $j$ at client $k$, and (ii) the aggregated privacy noise $\bm{M}_{k}^{t,-j}$. The challenges are threefold: (i) The privacy noises $\bm{N}_{m}^{t}$ are correlated, $\nu_{k}^{t}$ and $\nu_{k}^{t,-j}$ do not vary monotonically with the number of participating devices; (ii) The $\nu_{k}^{t}$ and $\nu_{k}^{t,-j}$ are random variables with respect to the network topology; (iii) The message to be protected, i.e., the $j$-th contribution $g_{k,j}(\Delta\bm{\Theta}_{j}^{t}+\bm{N}_{j}^{t})$, is unbounded. These challenges are addressed through the choice of the Bernstein parameter [4] and the Chi-squared tail bound.
Theorem V.2 (LDP in Protecting Identity in Partial Sums).
@@
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.3, then the variance of $\bm{M}_{k}^{t}$ is given by $\nu_{k}^{t,-j}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t,-j}\bm{A}\rVert^{2}\bm{I}_{D}$, and $\bar{\nu}_{k}^{-j}=\mathbb{E}[\nu_{k}^{t,-j}]=\left(\lambda^{2}\sum_{m=1}^{K}(1-p_{k,m})g_{m,k}^{2}+\bm{\Lambda}_{k}^{\mathrm{E},-j}\bm{A}\bm{A}^{\top}(\bm{\Lambda}_{k}^{\mathrm{E},-j})^{\top}\right)$, where $\bm{\Lambda}_{k}^{\mathrm{E},-j}=\mathbb{E}[\bm{\Lambda}_{k}^{t,-j}]$. For $\forall\delta_{0}$, $r_{1}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{k}^{t,-j}-\bar{\nu}_{k}^{-j}\rvert\geq r_{1})\leq\delta^{\prime}$, [^12] The choice of Bernstein parameter $r_{1}$$r_{2}$ $r_{3}$$r_{4}$ is discussed in Appendix F. the partial sum $\bm{S}_{k}^{t}$ is $(\epsilon_{k,j}^{(2)},(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(2)}))$-differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any participating node $j$, where
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t,-j}\in\mathcal{S})\leq e^{\epsilon_{k,j}^{(2)}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S})+(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(2)}),$ (29)
@@
Apart from the participating identity, when the transmission from client $j$ is perturbed from $\Delta\bm{\Theta}_{j}^{t}=\bm{u}$ to $\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime}$, the aggregated partial sums should not fluctuate significantly. Different from the previous case, the noise offering privacy protection here is $\bm{M}_{k}^{t}$, which includes the $j$-th secret key.
Theorem V.3 (LDP Against Perturbation of Local Data in Partial Sums).
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.3, then its variance is $\nu_{k}^{t}\bm{I}_{D}=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}$. The $\bar{\nu}_{k}=\mathbb{E}[\nu_{k}^{t}]$ and $\bm{\Lambda}_{k}^{\mathrm{E}}=\mathbb{E}[\bm{\Lambda}_{k}^{t}]$ are specified in Theorem V.2. For $r_{2}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{k}-\bar{\nu}_{k}\rvert\geq r_{2})\leq\delta^{\prime}$, the partial sum $\bm{S}_{k}^{t}$ is $(\epsilon_{k,j}^{(3)},(1-p_{k,j})(\delta^{\prime}+\delta_{j}^{(3)})$-differentially private in protecting the local date of any participating node $j$ from perturbations, where
$\epsilon_{k,j}^{(3)}=\begin{cases}2\cdot\left[2\log\left(\frac{1.25}{\delta_{j}^{(3)}}\right)\right]^{\frac{1}{2}}\frac{\lvert g_{k,j}\rvert R}{\sqrt{\bar{\nu}_{k}-r_{2}}}\;\;\;\;\;\mathrm{if}\;p_{k,j}\geq 0,\\ \hskip 42.67912pt\;\;\;\;\;0\;\;\;\;\;\;\hskip 42.67912pt\;\;\;\mathrm{if}\;p_{k,j}=0.\end{cases}$ (30)
@@
$\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u})$ $\leq e^{\epsilon_{k,j}^{(3)}}\mathrm{P_{r}}(\bm{S}_{k}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})$
@@ V-C LDP at the Server
Assume that the server receives partial sums from clients in $\mathcal{V}^{t}\triangleq\{k:\tau_{k}^{t}=1\}$, but fails to reconstruct the global model using the combinator $\bm{c}_{f^{\prime}}$ (referred to as a $\bm{c}_{f^{\prime}}$-Failure). The resulting aggregation result $\bm{F}_{f^{\prime}}^{t}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t,I}+\bm{N}_{m}^{t})$ still contains residual privacy noise. The effective noise in this aggregated result is given by $\bm{N}_{F_{f^{\prime}}}^{t}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$. The variance of $\bm{N}_{F_{f^{\prime}}}^{t}$ is $\nu_{F_{f^{\prime}}}^{t}\bm{I}_{D}=\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})$, and its expectation is $\bar{\nu}_{F_{f^{\prime}}}^{t}=\mathbb{E}[\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})]$. Similarly, define $\bm{F}_{f^{\prime}}^{t,\{-j,k\}}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t,I}+\bm{N}_{m}^{t})$ as the aggregation excluding client $j$’s contribution at relaying client $k$, and $\bm{N}_{F_{f^{\prime}}}^{t,\{k,-j\}}=\sum_{k\in\mathcal{V}^{t}}c_{f^{\prime},k}\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}$. Moreover, define $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}=\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t,\{k,-j\}})$ and $\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}}=\mathbb{E}[\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t,\{k,-j\}})]$ . The $\nu_{F_{f^{\prime}}}^{t}$, $\bar{\nu}_{F_{f^{\prime}}}^{t}$, $\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}$ and $\bar{\nu}_{F_{f^{\prime}}}^{\{k,-j\}}$ are further specified in Appendix G. Let $\mathcal{V}_{f^{\prime}}\triangleq\{k:c_{f^{\prime},k}\neq 0\}$, and $\mathcal{V}_{f^{\prime}}^{t}\triangleq\mathcal{V}_{f^{\prime}}\cap\mathcal{V}^{t}$.
From $\bm{F}_{f^{\prime}}^{t}$, the server should not infer the participation of client $j$ at the relaying client $k$. This privacy is offered by: (i) the privacy noise $\bm{N}_{F_{f^{\prime}}}^{t,\{k,-j\}}$, which consists of: (a) the privacy noises sent from other clients (excluding client $j$), $\mathcal{U}_{k}\setminus\{j\}$, to client $k$; and (b) the privacy noises sent from $\mathcal{U}_{m}$ to client $m\in\mathcal{V}_{f^{\prime}}\setminus\{k\}$, corresponding to non-zero combination coefficients $c_{f^{\prime},m}$; (ii) the link intermittency $\tau_{k,j}$ from client $j$ to client $k$; and (iii) the link intermittency $\tau_{k}$ from client $k$ to the server.
If $\Delta\bm{\Theta}_{j}^{t}=\bm{u}$ is perturbed to $\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime}$, the server should not observe a significant fluctuation in the resulting $\bm{F}_{f^{\prime}}^{t}$. The analytical difficulty arises as the perturbation propagates through multiple relaying nodes. The server can observe the perturbation unless every relaying path fails. The perturbation of $j$-th local data is protected by (i) the privacy noise $\bm{N}_{F_{f^{\prime}}}^{t}$, which consists of: (a) the privacy noises sent from clients (including client $j$), $\mathcal{U}_{k}$, to client $k$; and (b) the privacy noises sent from $\mathcal{U}_{m}$ to client $m\in\mathcal{V}_{f^{\prime}}\setminus\{k\}$, corresponding to non-zero combination coefficients $c_{f^{\prime},m}$; (ii) the link intermittency $\tau_{m,j}$ from client $j$ to all clients $m\in\mathcal{V}_{j}$ that relay the $j$-th message; and (iii) the link intermittency $\tau_{m}$ from client $m\in\mathcal{V}_{j}$ to the server.
Lemma V.1.
The server cannot observe the perturbation of the local data on client $j$ through $\bm{c}_{f^{\prime}}$-combination, only if all relaying clients that successfully hear from client $j$ and corresponds to a non-zero combination coefficient fails to communicate with the server, i.e., for $\forall m\in\mathcal{V}_{j,f^{\prime}}^{t}\triangleq\mathcal{V}_{j}^{t}\cap\mathcal{V}_{f^{\prime}}$, $\tau_{m}=0$, where $\mathcal{V}_{j}^{t}\triangleq\{m:m\in\mathcal{V}_{j}\}\cap\{m:\tau_{m,j}\neq 0\}$ denotes the set of neighboring clients who successfully hears from client $j$, and $\mathcal{V}_{f^{\prime}}\triangleq\{k:c_{f^{\prime},k}\neq 0\}$ denotes the clients who correspond to a non-zero combination coefficients in $\bm{c}_{f^{\prime}}$. As a result, the probability of this event is given by
$\tilde{p}_{j,f^{\prime}}=\hskip-2.84526pt\sum_{\mathcal{V}_{j,f^{\prime}}^{t}\in\{\mathcal{V}_{j}\cap\mathcal{V}_{f^{\prime}}\}}\prod_{k_{1}\in\mathcal{V}_{j,f^{\prime}}^{t}}(1-p_{k_{1}})\cdot p_{k_{1},j}\hskip-2.84526pt\hskip-8.53581pt\prod_{k_{2}\in\{\mathcal{V}_{j}\cap\mathcal{V}_{f^{\prime}}\}\setminus\mathcal{V}_{j,f^{\prime}}^{t}}\hskip-8.53581ptp_{k_{2}}.$ (32)
Theorem V.4 (LDP in $\bm{c}_{f^{\prime}}$-Failure).
@@
Given secret keys $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ generated by Construction IV.3 and the combinator $\bm{c}_{f^{\prime}}$, then for any $r_{3}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{F_{f^{\prime}}}^{t,\{k,-j\}}-\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}\rvert\geq r_{3})\leq\delta^{\prime}$, then, in $\bm{c}_{f^{\prime}}$-Failure, $\bm{F}_{f^{\prime}}^{t}$ is $(\epsilon_{f^{\prime},\{k,j\}}^{(4)},(1-p_{k})(1-p_{k,j})(\delta_{f^{\prime},\{k,j\}}^{(4)}+\delta^{\prime})$ differentially private w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$ in protecting the identity of any participating client $j$ at relay $k$, where
@@
$\epsilon_{f^{\prime},\{k,j\}}^{(4)}\hskip-2.84526pt=\hskip-2.84526pt\left[\hskip-1.42262pt2\log\left(\hskip-2.84526pt\frac{1.25}{\delta_{f^{\prime},\{k,j\}}^{(4)}}\hskip-2.84526pt\right)\hskip-4.2679pt\right]^{\frac{1}{2}}\hskip-5.69054pt\frac{\lvert c_{f^{\prime},k}g_{k,j}\rvert(R+\lambda\sqrt{D(1+\delta_{0})})}{\sqrt{\bar{\nu}_{F_{f^{\prime}}}^{t,\{k,-j\}}-r_{3}}}.$ (33)
@@
Moreover, for any $r_{4}>0$ and $\delta^{\prime}\in(0,1)$ such that $\mathrm{P_{r}}(\lvert\nu_{F_{f^{\prime}}}^{t}-\bar{\nu}_{F_{f^{\prime}}}^{t}\rvert\geq r_{4})\leq\delta^{\prime}$, for any $\bm{u},\bm{u^{\prime}}\in\mathbb{B}_{D}(R)$, it holds that
@@
$\epsilon_{k,j,f^{\prime}}^{(5)}=2\left[2\log\left(\frac{1.25}{\delta_{j}^{(5)}}\right)\right]^{\frac{1}{2}}\frac{\sum_{k=1}^{K}\lvert c_{f^{\prime},k}g_{k,j}\rvert R}{\sqrt{\bar{\nu}^{t,f^{\prime}}-r_{4}}}.$ (36)
@@
With a slight abuse of the paper organization, we examine privacy via obfuscation by other local models in successful server aggregation, considering the stochasticity of the local models. Unlike the previous LDP analysis, where noise is explicitly added to each input, privacy arises naturally from the randomness or uncertainty in the other local models. For this reason, the server cannot see significant changes in the global model $\bm{M}^{t}=\frac{1}{K}\sum_{k=1}^{K}\Delta\bm{\Theta}_{k}^{t}$ resulting from the presence or perturbation of $\Delta\bm{\Theta}_{j}^{t}$. To facilitate theoretical analysis, it is assumed $\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])=\mathrm{Diag}(\bm{I}_{D}\zeta^{2})$ in this analysis.
Theorem V.5 (LDP in Global Model).
@@
w.p. at least $1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}$. Moreover, for any $\bm{u},\bm{u^{\prime}}$, it holds that
@@
$\leq e^{\epsilon^{(7)}}\mathrm{P_{r}}(\bm{M}^{t}\in\mathcal{S}|\Delta\bm{\Theta}_{j}^{t}=\bm{u}^{\prime})+\delta^{(7)}.$ (39)
w.p. at least $\left(1-e^{-\frac{D}{2}(\delta_{0}-\ln{(1+\delta_{0})})}\right)^{2}$, where
@@
Remark V.1.
If $K\rightarrow\infty$, then $\epsilon^{(6)}\rightarrow 0$ and $\epsilon^{(7)}\rightarrow 0$.
VI Convergence
A non-convex convergence rate analysis of SecCoGC under an arbitrary choice of local solver is conducted. First, the assumptions concerning LOFs are present. The combination of assumptions VI.1VI.3 is the most relaxed in the non-convex FL analysis [21] and allows for broad applicability in real-world settings involving deep neural networks (DNNs).
Assumption VI.1 (Smoothness).